OpenStudy (anonymous):

Trying to find the x and y intercepts of x=3y^2-7

OpenStudy (anonymous):

Set x equal to 0 to solve for the y intercept. Set y equal to 0 for the x-intercept. Can you take it from there?

OpenStudy (anonymous):

I set y = to 0 and get x=-4 so the X intercept is (0,-4)?

OpenStudy (anonymous):

How exactly did you get -4? Can you show me what you substituted?

OpenStudy (anonymous):

x=3(0)^2-7 ah but its -7 so X intercept is (-7,0)?

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

o know the Y intercept is the one I am really having trouble with because when working it I I have the square and a fraction

OpenStudy (anonymous):

Okay. x=3y^2-7 0=3y^2-7 7=3y^2 7/3=y^2 \[\pm \sqrt{\frac{ 7 }{ 3 }}\]

OpenStudy (anonymous):

Do you need me to explain the process, or are you fine with that?

OpenStudy (anonymous):

I did get that. I typed it in before had into my homework and got it wrong. I typed: (0,-7/3)(0,7/3)

OpenStudy (anonymous):

huh. No clue.

OpenStudy (anonymous):

Have you tried reversing the answers (negative before the positive coordinates)?

OpenStudy (anonymous):

yes